Construct a three-step synthesis of trans-2-pentene from acetylene by dragging the appropriate formulas into the bins. Note that each bin will hold only one item, and not all of the given reagents or structures will be used. (Me = methyl, CH3; Et = ethyl, CH3CH2.)

Question 20 of 20 sapling learning Construct a three-step synthesis of trans-2-pentene from acetylene by dragging the appropriate formulas into the bins. Note that each bin will hold only one item, and not all of the given reagents or structures will be used. (Me methyl, CH3, Et ethyl, CH3CH2.) > · Product (trans-2-pentene) Reactant Reagent 1 1Stap 1 Step 1 Product Reagent 2 Step 2 Product Reagent 3 (acetylene) Na HBr H2 NH3 H 1) NaNH2 H 2) MeBr 1) NaNH2 2) EtBr Previous Give Up & View Solution O Check Answer ) Next Exit Hint
Answer #1
Concepts and reason

The concept of this question is conversion of an alkyne to an alkene.

Alkynes being unsaturated undergo addition reactions. The important chemical reactions of alkynes include electrophilic addition reactions, nucleophilic addition reactions , reactions due to acetylenic hydrogen, polymerisation , isomerization and oxidation reactions.

Fundamentals

The conversion of an alkyne to alkene takes place in generally two steps which includes the preparation of higher alkynes from lower alkynes followed by the dissolved metal reduction of the higher order alkyne in order to make an alkene.

The reaction for conversion of acetylene (ethyne) to one higher order alkyne is as follows:

HC=CH NINH> HC=C NACHBT-HC=C-CH,
(Higher order Alkyne)
+
NaBr

The reaction is as follows:

CH,CH,Br
HC=C-CH, NaNH> NaC=C-CH CÓ
NaBr + CH.-H.C-C=C-CH
CH.HCC.ch
NaBr
(Higher order Alkyne)

The reaction is as follows:

HC-H.C.
CH,-HỌC-C=C-CH, Na, >
CEC
H
Dissolving mutal
reduction
Trans-Pent-2-ene

Ans:

The correct reagents and products in overall reaction is as follows:

HCSCH CAB:> HC=C-CH, CH,CH;B[7
NaNH,
HỌC-HC
- UI1112
C=C
H, Na CH-HTC-C =C-CH,
Pent-2-yne
SCH;
NH,
Trans-Pent-2-ene